OpenStudy (anonymous):

Solving Equations with Radicals. (sqrt of x)+1=5

OpenStudy (anonymous):

subtract 1 from both sides so it is x^2=4 then find the square root of both sides

OpenStudy (demandandsupply):

* then square both sides

OpenStudy (anonymous):

(sqrt of x)+1=5 x+1 x-1=5-1 x=4 like that?@mathisfrustrating @DemandAndSupply

OpenStudy (anonymous):

x=2, because the square root of 4 = 2